OpenStudy (anonymous):

simplify: [(n+1)!/n!]-n

OpenStudy (zarkon):

\[(n+1)!=(n+1)\times n!\]

OpenStudy (anonymous):

(n+1!/(n-1)! = (n+1)*n*(n-1)(n-2).../(n-1)*(n-2)... = (n+1)n OR n(n+1) Whenever doing division problems with factorials,always look for the results of cancelation of terms in the numerator by those in the denominator.

OpenStudy (anonymous):

does that help:)

OpenStudy (anonymous):

still a bit confused :/